VIBRATIONAL POTENTIAL ENERGY SURFACES FOR THE RING-PUCKERING AND RING-FLAPPING VIBRATIONS OF 1,3-BENZODIOXOLE IN ITS S_{0} AND S_{1} (π, π^{∗}) ELECTRONIC STATES

\underline{J. LAANE}, K. MORRIS, S. SAKURAI, E. BONDOC, Department of Chemistry, Texas A&M University, College Station, TX 77842-3012; N. MEINANDER, Department of Physics, University of Helsinki, Helsinki, Finland.